OpenStudy (anonymous):

I now just need help finding the upper range of the below equation. I have the domain as (-3,3) and the range [2/(sqrrt(3), ?????] thanks for the help!! Please help me find the domain and range of : y=2/(4th root(9-x^2)

OpenStudy (anonymous):

Domain is first number and range would be the second number. so just find the x and y

OpenStudy (rsadhvika):

which part are you stuck at

OpenStudy (freckles):

having even roots in a problem you will have some restrictions if the thing inside is less than 0 also having fractions can be a problem because we don't want to divide by 0 you need to make sure the thing inside your 4th root there is positive

OpenStudy (anonymous):

so is the domain -3 to 3?

OpenStudy (anonymous):

but not including 3?

OpenStudy (freckles):

sounds good not including either -3 or 3

OpenStudy (anonymous):

so then do you put -3 and 3 into the equation to get the range limits?

OpenStudy (anonymous):

but then that would give you zeros for the denominator so what then?

OpenStudy (freckles):

what happens when you take 1 and divide it by numbers that get really really close to 0?

OpenStudy (freckles):

like what is 1/.00001 equal to? is that actually a big number?

OpenStudy (anonymous):

yeah infinity

OpenStudy (freckles):

and we restricted our domain so that the function exists over the real numbers we basically have a positive divided by a positive so we know that are range will exist only for positive numbers and it actually goes to infinity but we also need to figure out the lowest y value

OpenStudy (freckles):

the thing will be the smallest if you replace x with 0 and solve for y

OpenStudy (anonymous):

that would be 2/(4th rt(9) - our teacher makes us do specific answers, no decimals

OpenStudy (anonymous):

so that is the low end of our range

OpenStudy (freckles):

you could simplify that though

OpenStudy (anonymous):

so the top end would be when x is close to 3 but not 3?

OpenStudy (freckles):

\[\sqrt[4]{9}=\sqrt[4]{3^2}=3^\frac{2}{4}=?\]

OpenStudy (anonymous):

sorry my computer kicked me off

OpenStudy (anonymous):

I am not sure, I see how you got to that point though

OpenStudy (anonymous):

oh duh, are you just saying 3^1/2?

OpenStudy (anonymous):

so 2/sqrt(3)?

OpenStudy (freckles):

well yeah that would be the smallest of the range

OpenStudy (freckles):

and then you noticed that if you plug in either endpoint the bottom was getting closer to zero therefore the fraction was getting positive large

OpenStudy (freckles):

so do you what the range is?

OpenStudy (freckles):

also you weren't the only one having problems i couldn't get openstudy to load for like 5 to 6 minutes

OpenStudy (anonymous):

Any help on the upper part of the range? I get that x=0 at the low end, and x should = 3 at the upper end, but 3 cannot be included, so it is a number infanetly close to three and i can't wrap my brain around that?

OpenStudy (freckles):

emily we already got the upper part

OpenStudy (anonymous):

and the domain doesn't include -3 or 3, but would the range include both numbers? I feel like it should include them?

OpenStudy (freckles):

remember you said the the function was getting huge as our x's approaches x=3 or x=-3

OpenStudy (freckles):

the domain is (-3,3)

OpenStudy (freckles):

the range is [2/sqrt(3),inf) because the smallest was 2/sqrt(3) and because the function got huge when the bottom went to 0

OpenStudy (freckles):

got positive huge*

OpenStudy (freckles):

emily do you understand?

OpenStudy (anonymous):

Oh wow, yeah duh, it would be infinity wouldnt it? Thanks!